A sample of 52 copper wires had a mean resistance of 1.03 mΩ. (Assume that the standard deviation for resistance is known to be 0.1 mΩ.) Let μ represent the mean resistance of copper wires of this type.

Find the P-value for testing H0 : μ ≤ 1 versus H1 : μ > 1.
Answer #1

this is the right tailed test .  

The null and alternative hypothesis is ,

H0 : μ ≤ 1 versus H1 : μ > 1

Test statistic = z

= ( - ) / / n

= (1.03 - 1) / 0.1/ 52

= 2.16

P(z > 2.16) = 1 - P(z < 2.16) = 1 - 0.9846=0.0154

P-value = 0.0154